Chaundra is two years older then her brother. She is four years younger then her older sister. Her two parents are each forty-one. The total of all five ages is one hundred twenty-three. How old is Chaundra?
Answer by WannabeCAgirl83(35)

Chaundra = x
Chaundra's younger brother = x - 2
Chaundra's older sister = x + 4
Chaundra's mother = 41
Chaundra's father = 41
all 5 of them together = 123

Chaundra + brother + sister + mother + father = 123

x + x-2 + x+4 + 41 + 41 = 123 | re-arranging
3x + 84 = 123 | - 84 on both sides
3x = 39 | ÷ 3 on both sides
x = 13 → Chaundra's age

You can substitute x by 13 to check if your solution is correct (if the result really is 123):

13 + 13-2 + 13+4 + 41 + 41 = 123 → true